# Multi-modal deformable image registration using untrained neural networks

### v1

@@ Abstract
Image registration techniques usually assume that the images to be registered are of a certain type (e.g. single- vs. multi-modal, 2D vs. 3D, rigid vs. deformable) and there lacks a general method that can work for data under all conditions. We propose a registration method that utilizes neural networks for image representation. Our method uses untrained networks with limited representation capacity as an implicit prior to guide for a good registration. Unlike previous approaches that are specialized for specific data types, our method handles both rigid and non-rigid, as well as single- and multi-modal registration, without requiring changes to the model or objective function. We have performed a comprehensive evaluation study using a variety of datasets and demonstrated promising performance.
@@ 1 Introduction
Fig. 1: Untrained neural network for image registration. (a) The model outputs two channels, each representing a different modality, and two time points, denoting fixed and transformed input images. The L2 loss between the reconstructed images and their corresponding ground truth images is then computed and used to update the networks’ weights. For single-modal registration, the L2 loss is applied on one output channel while in the multi-modal case, the loss function is computed using different output channels corresponding to the image modalities. The hash embedding is always applied to the input coordinate, so we consider it part of the network and omit it for visibility. (b) The model can handle rigid and deformable registration by tuning the granularity value.
Recently, neural networks trained on datasets with paired images and registration labels have demonstrated success in handling large deformations [6]. Nonetheless, curating a large dataset with ground truth labels is often infeasible, particularly for multi-modal images. In the absence of such training dataset, an image-to-image network have been used to convert an image from one modality to the other, simplifying the problem to single-modal registration [7, 8]. Alternatively, untrained coordinate-based neural networks are shown to be a good building block for image registration methods, providing a flexible way to parameterize the spatial transformation between two images [9, 10]. Byra et al. [10] has used such network to register single-modal 2D brain images. In our work, we extend to multi-modal deformable registration by combining two coordinate-based networks.
@@ 2.2 Image registration using untrained neural networks
The image network outputs one channel for single-modal and two channels multi-modal registration. For the multi-modal case, the first and second terms of the loss function are applied on different output channels, one for each image modality.
@@ 3.1 Datasets and metrics
For deformable registration, we used the Abdomen MR-CT dataset [14, 15], which consists of 16 pairs of MR-CT scans with manual segmentations of different organs and arbitrary deformable motion between the images. Since our method does not require training, we only use the first six pairs of MR-CT scans from the Abdomen MR-CT dataset for our evaluation. For quantitative evaluation, we use dice similarity coefficient (DSC) which measures the volumetric overlap between the ground truth segmentation and the estimated segmentation, and HD-95 which assesses $95^{\text{th}}$ percentile of the distances between the points of the ground truth and the predicted labels.
@@ 3.3 Multi-Modal Registration
Average DSC $0.43\pm 0.17$ $0.47\pm 0.20$ $\mathbf{0.51\pm 0.11}$
Weighted DSC $0.53\pm 0.16$ $0.55\pm 0.13$ $\mathbf{0.56\pm 0.14}$
HD-95 (mm) $18.3\pm 8.50$ $\mathbf{17.9\pm 8.87}$ $18.9\pm 8.86$
For 3D multi-modal and deformable registration, we use symmetric normalization SyN (affine and deformable transformation) with mutual information (MI) and cross-correlation (CC) as the optimization metric as the baseline methods [17]. As shown in Table 3, our proposed method achieves a higher average and weighted Dice than both settings of ANTs. However, it receives a slightly worse HD-95, $1.0$ mm higher, than the baseline method that uses CC.
Our model is able to register multi-modal images without explicitly optimizing for the pixel correlation. As illustrated in Fig. 2, without restricting the network capacity in the earlier training, the image MLP is prone to overfit to both input images too quick before the motion MLP can account for any motion. However, with the coarse-to-fine process, the model capacity is limited in the beginning, forcing the motion network to search for the best alignment and resulting in a good registration.
@@ 4 Discussion
Our findings are subjected to several challenges. Since the implicit networks are trained from scratch for each new image pair, their inference time is slower compared to conventional convolutional networks. It would be interesting to investigate how hyper-networks can be utilized to pre-initialize the weights of the networks or how MLPs can be pre-trained on population data [18]. Furthermore, our current use of a simple L2 loss without regularization may result in non-smooth and non-invertible deformation fields [19]. Incorporating more sophisticated regularizers for weights optimization could further enable the model to better capture larger and more complex deformations.

### v2

@@ Abstract
Image registration techniques usually assume that the images to be registered are of a certain type (e.g. single- vs. multi-modal, 2D vs. 3D, rigid vs. deformable) and there lacks a general method that can work for data under all conditions. We propose a registration method that utilizes neural networks for image representation. Our method uses untrained networks with limited representation capacity as an implicit prior to guide for a good registration. Unlike previous approaches that are specialized for specific data types, our method handles both rigid and non-rigid, as well as single- and multi-modal registration, without requiring changes to the model or objective function. We have performed a comprehensive evaluation study using a variety of datasets and demonstrated promising performance.  [^1] Code: https://github.com/quang-nguyenln/mdirunn.
@@ 1 Introduction
Fig. 1: Untrained neural network for image registration. (a) The model outputs two channels, each representing a different modality, and two time points, denoting fixed and transformed input images. The L2 loss between the reconstructed images and their corresponding ground truth images is then computed and used to update the networks’ weights. For single-modal registration, the L2 loss is applied on one output channel while in the multi-modal case, the loss function is computed using different output channels corresponding to the image modalities. The hash embedding is always applied to the input coordinate, so we consider it part of the network and omit it for visibility. (b) The model can handle rigid and deformable registration by tuning the granularity value. (c) An example of the 3D multi-modal registration of MRI-CT data. The green contours are the liver and spleen segmentations annotated on the displayed image, and the red contour refers to the segmentation annotated on the other image modality and warped using our method.
Recently, neural networks trained on datasets with paired images and registration labels have demonstrated success in handling large deformations [6]. Nonetheless, curating a large dataset with ground truth labels is often infeasible, particularly for multi-modal images. In the absence of such training dataset, an image-to-image network have been used to convert an image from one modality to the other, simplifying the problem to single-modal registration [7, 8]. Alternatively, untrained coordinate-based neural networks are shown to be a good building block for image registration methods, providing a flexible way to parameterize the spatial transformation between two images [9]. Byra et al. [9] has used such network to register single-modal 2D brain images. In our work, we extend to multi-modal deformable registration by combining two coordinate-based networks.
@@ 2.2 Image registration using untrained neural networks
The image network outputs one channel for single-modal and two channels for multi-modal registration. For the multi-modal case, the first and second terms of the loss function are applied on different output channels, one for each image modality.
@@ 3.1 Datasets and metrics
For deformable registration, we used the Abdomen MR-CT dataset [13, 14], which consists of 16 pairs of MR-CT scans with manual segmentations of different organs and arbitrary deformable motion between the images. We use the eight pairs of MR-CT scans from the Abdomen MR-CT dataset that have ground truth labels for our evaluation. For quantitative evaluation, we use dice similarity coefficient (DSC) which measures the volumetric overlap between the ground truth segmentation and the estimated segmentation, and HD-95 which assesses $95^{\text{th}}$ percentile of the distances between the points of the ground truth and the predicted labels.
@@ 3.3 Multi-Modal Registration
Average DSC $0.41\pm 0.21$ $0.44\pm 0.23$ $\mathbf{0.49\pm 0.17}$
Weighted DSC $0.51\pm 0.15$ $0.52\pm 0.14$ $\mathbf{0.54\pm 0.14}$
HD-95 (mm) $19.6\pm 11.1$ $\mathbf{18.5\pm 11.2}$ $18.9\pm 10.1$
Fig. 3: The network reconstruction of fixed image after applying motion kernel (green channel) overlaid with the transformed image (red). The left image is registered using an image network with full representation capacity, while the right image is registered with an image MLP with reduced capacity.
For 3D multi-modal and deformable registration, we use symmetric normalization SyN (affine and deformable transformation) with mutual information (MI) and cross-correlation (CC) as the optimization metric as the baseline methods [16]. As shown in Table 3, our proposed method achieves a higher average and weighted Dice than both settings of ANTs. However, it receives a slightly worse HD-95, $0.4$ mm higher, than the baseline method that uses CC.
Our model is able to register multi-modal images without explicitly optimizing for the pixel correlation. As illustrated in Fig. 2, without restricting the network capacity in the earlier training, the image MLP is prone to overfit to both input images too quickly before the motion MLP can account for any motion. However, with the coarse-to-fine process, the model capacity is limited in the beginning, forcing the motion network to search for the best alignment and resulting in a good registration.
@@ 4 Discussion
We simulate the extreme situation where the image contains more information than what the image network can store by limiting its granularity to $\frac{1}{15}$ the original resolution. Fig. 3 shows that our method maintains robust registration performance in such situation.
Our findings are subjected to several challenges. Since the implicit networks are optimized from scratch for each new image pair, their runtime is longer than convolutional networks trained with paired data. It would be interesting to investigate how hyper-networks can be utilized to pre-initialize the weights of the networks or how MLPs can be pre-trained on population data [17]. Furthermore, incorporating more sophisticated regularizers for weights optimization could prevent convergence to non-smooth and non-invertible deformation fields [18] and enable the model to better capture larger and more complex deformations.